A newspaper earns a profit of $20 per year for 3000 subscribers. They project that the profit will increase by .01 for each additional subscriber...how many subscribers are needed to bring the total profits to $138125.....
Answer by LinnW(1048)

we want the equation of a line with slope = 0.01 through (3000,20)
using y = mx + b as a template,
20 = (0.01)(3000) + b
20 = 30 + b
add -30 to each side
-10 = b
So our equation is y = (0.01)x -10
The number of subscribers x for $138125 is
138125 = (0.01)x -10
add 10 to each side
138115 = 0.01x
multiply each side by 100
13811500 = x
So we need 138,811,500 subscribers
Checking our equation for 3000 subscribers,
y = 0.01(3000) - 10
y = 30 - 10
y = 20